### v2

@@ B Mathematical Proofs
where we ignore the subscript in (14) and denote $\widetilde{\mathbf{F}}^{s}$ as the estimated factor matrix from $\widetilde{\mathbf{X}}_{1}$ with sparsity $s$, and ${\mathbf{F}}^{s}$ as the true factor matrix retaining only keeps the top $s$ elements, both of which are normalized.
Proof. We only show it for $r=1$ since the proof is similar for $r>1$. Suppose the true factor ${\mathbf{F}}^{s}$ is the normalized true factor matrix such that ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$, which is similarly assumed in Bai and Ng (2013). By (14), we have
@@
Let Assumptions 17 hold. For any $s$ with $s<s_{0}$, there exists a constant $\tau_{s}>0$ such that
@@
Since each column in ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}$ is a hat matrix and ${\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}$ is a positive semi-definite matrix, then there exists a constant $\tau_{s}>0$ such that $I\geq\tau_{s}N_{2}T$. Then Lemma A.9 follows from this result and (S.24)–(B). This completes the proof. $\Box$
@@
Let Assumptions 17 hold. For any $s$ with $s\geq s_{0}$, we have
@@
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}$ (S.31)
@@
By a similar argument for the remaining terms in (B), we can show that
@@
Proof of Theorem 4. We only consider the case when the number of partitions $J=1$ since it is similar for a genera $J>1$. We will prove that
$\lim_{N,T\rightarrow\infty}P(PC(s)<PC(s_{0}))=0,$
for all $s\neq s_{0}$ and $s\leq smax$. We first consider the case when $s<s_{0}$. Note that
$PC(s)-PC(s_{0})=R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})-r\frac{(s_{0}-s)}{T^{\alpha}}g(N_{1},T),$
@@
$P(PC(s)-PC(s_{0})<0)=P(\tau_{s}+O_{p}(C_{N_{1}T})<r\frac{(s_{0}-s)}{T^{\alpha}}g(N_{1},T))\rightarrow 0,$
since $g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}\rightarrow 0$.
@@
$P(PC(s)-PC(s_{0})<0)=P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r\frac{(s-s_{0})}{T^{\alpha}}g(N_{1},T)).$
@@
Since $C_{N_{1}T}^{-1}g(N_{1},T)\rightarrow\infty$, we have that
$P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r(s-s_{0})g(N_{1},T))\rightarrow 0,$
as $N,T\rightarrow\infty$. This completes the proof. $\Box$
C Description of the Identified Sparse Factors
@@
Table S.I: Significant sparse factors over the time horizon from January 1, 2004, to December 31, 2016. Dates are formatted as yyyymmdd. The signs in the Factor Return column indicate whether the significant time factor resulted in a positive ($+$) or negative (--) return on that day. Reasons in the Reason column are extracted from daily reports on CNN Money (www.money.cnn.com) after the market closes each trading day. The Time Factor column summarizes the corresponding factor for each trading day with significant common risk.
@@
20070227 -- A big decline in Chinese stocks, weakness in some key readings on the US economy and news that Vice President Dick Cheney was the apparent target in a Taliban suicide bombing attack in Afghanistan China, Economic indicators, Global events (terrorist attack)
@@
20070803 -- Credit market fears, sparked by Wall Street bank Bear Stearns Credit risk, Market sentiment
@@
20070918 $+$ After the Federal Reserve cut a key short-term interest rate by a half-percentage point Government policies
20071107 -- Renewed credit market fears Credit risk, Market sentiment
@@
20071128 $+$ Expectations the Federal Reserve will continue its rate cutting campaign and on strength from the embattled financials Government policies, Market sentiment
@@
20071211 -- After the Federal Reserve cut the fed funds rate by a quarter-percentage point, as expected, but disappointed some investors looking for a bigger cut Government policies, Market sentiment
@@
20080104 -- After a weaker-than-expected December jobs report exacerbated worries that the economy may be falling into recession Economic indicators, Market sentiment
@@
20080117 -- Recession worries following comments from Federal Reserve Chairman Ben Bernanke, Merrill Lynch’s big quarterly loss and weak readings on the housing and manufacturing sectors Market sentiment, Company-specific factors, Economic indicators
@@
20080123 $+$ The Federal Reserve stepped in Tuesday and announced an emergency intermeeting interest rate cut, a decision that initially had a mixed impact on stocks, but helped pave the way for Wednesday’s bounce back Government policies
@@
20080205 -- After a report showing a big slowdown in the services sector of the economy and cautionary comments from a Fed official amplified fears that a recession is underway or imminent Company-specific factors, Government policies
@@
20080229 -- After AIG’s record loss added to worries about the financial sector and more weak economic news intensified fears about a recession Company-specific factors
@@
20080311 $+$ Announcement: the Federal Reserve will lend up to \$$200$ billion to banks and lenders as a means of loosening up tight credit markets Government policies
@@
20080318 $+$ After the Federal Reserve cut the fed funds rate by three-quarters of a percentage point, surprising investors looking for a larger cut Government policies
@@
20080401 $+$ Investors cheered signs that the companies hit hardest by the credit market crisis seem to be working through the problems Market sentiment
@@
20080606 -- After oil prices spiked more than \$currency-dollar\\$\$$11$ a barrel and the May jobs report showed a big jump in the unemployment rate Oil price, Economic indicators
20080626 -- Selling accelerated following a record surge in oil prices Oil price
20080708 $+$ Falling oil prices and a stronger dollar Oil price
@@
20080709 -- More worries about Freddie Mac and Fannie Mae’s ability to raise capital exacerbated credit market and corporate profit jitters Market sentiment
@@
20080716 $+$ Encouraging news from the banking and airline sectors Falling oil prices also helped spark a strong stock market rally Economic indicators, Oil price
@@
20080717 $+$ A string of large oil price declines and more encouraging earnings results from the financial sector Oil price, Economic indicators
@@
20080724 -- Renewed fears about the battered housing market and rising unemployment Market sentiment
@@
20080729 $+$ Strong financial reports, rising consumer confidence and falling oil prices Economic indicators, Oil price
@@
20080805 $+$ Oil prices fell sharply and investors appeared to take solace in the Federal Reserve’s assessment of the nation’s economy Oil price, Government policies
@@
20080904 -- Mixed retail sales, lower oil prices and dour labor market readings amplified worries about a global economic slowdown Economic indicators, Oil price
@@
20080909 -- Worries about Lehman Brothers’ ability to raise capital, and about the extent of AIG’s mortgage-related losses, exacerbated broad recession fears Company-specific factors
@@
20080915 -- After one of the most calamitous days in US financial services history resulted in Bank of America’s \$$50$ billion deal to buy Merrill Lynch and the bankruptcy filing of Lehman Brothers Company-specific factors
@@
20080917 -- The government’s emergency rescue of AIG amplified fears about the stability of financial markets Government policies
@@
20080918 $+$ On a CNBC report that the government is working on a more permanent solution to absorbing bad debt Government policies
@@
20080919 $+$ The government’s plan to help rescue banks from toxic mortgage debt soothed investors at the end of a gut-churning week on Wall Street Government policies
@@
20080922 -- As investors worried about the specifics of the government’s \$$700$ billion bailout plan and rocketing oil prices - which saw its biggest one-day dollar gain ever Market sentiment
@@
20080929 -- After the House rejected the government’s \$$700$ billion bank bailout plan Government policies
@@
20080930 $+$ On bets that Congress will pass a version of the government’s \$$700$ billion package, following Monday’s crushing defeat Market sentiment, Government policies
@@
20081002 -- As frozen credit markets and weak economic reports amplified jitters ahead of the House vote on the \$$700$ billion bank rescue plan Market sentiment, Government policies
@@
20081006 -- As the \$$700$ billion bank bailout plan and European government attempts to prop up faltering banks failed to comfort panicky investors Market sentiment, Government policies
@@
20081007 -- As the Federal Reserve’s plan to loosen credit markets failed to temper investor pessimism Market sentiment, Government policies
@@
20081009 -- As panicked investors dumped stocks across the board Market sentiment
@@
20081013 $+$ As investors bet that the worst of the credit crisis is over, following a series of global initiatives announced over the last few days Market sentiment
@@
20081015 -- A weak retail sales report and dour forecasts from the Federal Reserve, coupled with sober comments from Fed Chairman Ben Bernanke, sent stocks tumbling Economic indicators, Government policies
@@
20081016 $+$ As the lowest oil prices in more than a year gave investors a reason to scoop up shares battered in the recent market selloff Oil price
@@
20081020 $+$ As investors welcomed talk of a second economic stimulus plan and an improvement in key lending rates Market sentiment, Government policies
@@
20081021 -- As mixed corporate earnings reports gave investors a reason to retreat after the previous session’s big rally Economic indicators
@@
20081022 -- As weak earnings and slumping oil prices amplified fears of a global recession Oil price
@@
20081024 -- As Wall Street joined a worldwide market slump on bets that a recession is imminent - if not already under way Market sentiment
@@
20081027 -- As recession jitters outweighed relief that the government’s programs to shore up the financial system have gotten underway Market sentiment
@@
20081028 $+$ As investors dove back into stocks near the end of one of the worst months in Wall Street history Market sentiment
@@
20081031 $+$ Capping off a strong week at the end of one of the worst months in Wall Street history Market sentiment
@@
20081104 $+$ As millions of Americans battered by the weakened economy turned out to vote for the next President of the United States Market sentiment
@@
20081105 -- As Barack Obama’s historic victory gave way to renewed worries about the struggling economy Market sentiment
@@
20081106 -- As fears of a prolonged recession sent investors running for the exits Market sentiment
@@
20081110 -- As ongoing recession fears overshadowed any relief about China’s \$$586$ billion stimulus plan and the government’s revamping of its deal with AIG China, Government policies
@@
20081111 -- As recession fears trumped a new government and mortgage industry plan to help troubled homeowners Market sentiment, Government policies
@@
20081112 -- As investors bet that a long and deep recession is on the horizon Market sentiment
@@
20081113 $+$ As the major stock gauges bounced back from levels not seen since 2003 Market sentiment
@@
20081114 -- As the worst retail sales on record ignited fears of a long recession Economic indicators
@@
20081117 -- As investors eyed Citigroup’s massive job losses and a weak manufacturing report, while awaiting the fate of a potential bailout for the automakers Company-specific factors, Economic indicators
@@
20081119 -- As ongoing anxiety about the economy and uncertainty about the future of the auto industry weighed on the market Market sentiment
@@
20081120 -- As fears of a prolonged recession sparked a massive selloff Market sentiment
@@
20081121 $+$ After reports surfaced that President-elect Barack Obama will nominate New York Federal Bank President Timothy Geithner as his new Treasury secretary Government policies
@@
20081124 $+$ As Citigroup’s massive rescue package and President-elect Obama’s picks for his economic team pushed investors off the sidelines Company-specific factors, Government policies
@@
20081126 $+$ As investors scooped up stocks hit in the recent selloff, ahead of the Thanksgiving holiday Market sentiment
@@
20081201 -- As investors bailed out following confirmation that the US is mired in a recession and indications that it’s likely to continue for some time Market sentiment
@@
20081202 $+$ As investors welcomed signs that the automakers might get a bailout after all Government policies
@@
20081204 -- As a rash of job cuts at major companies added to jitters ahead of the November jobs report Economic indicators
@@
20081205 $+$ After a brutal November employment report, as investors extended the recent trend of buying despite the bad news Economic indicators
@@
20081208 $+$ As investors welcomed President-elect Barack Obama’s plan to create jobs and revive the economy, and reports that government help for the automakers is on the way Government policies
@@
20081211 -- On worries that the \$currency-dollar\\$\$$14$ billion auto rescue bill won’t pass in the Senate due to Republican opposition Government policies
@@
20081216 $+$ After the Federal Reserve cut a key short-term interest rate to the lowest level on record, and signaled it had more tools available to help the economy as the recession stretches on Government policies, Market sentiment
@@
$2$- Amid concerns about fourth-quarter corporate earnings, falling oil prices and ongoing woes in the auto industry Market sentiment, Oil price, Economic indicators
@@
20081230 $+$ As investors scooped up a variety of shares hit hard in the 2008 stock market battering Market sentiment
@@
20090102 $+$ With investors starting off a new year on the right foot, after an abysmal 2008 Market sentiment
@@
20090106 $+$ As investors looked beyond the Federal Reserve’s dour outlook on the economy and instead scooped up shares hit in last year’s big selloff Market sentiment
@@
20090107 -- After weak labor market reports and dour forecasts from Alcoa and Intel gave investors reasons to retreat after the recent rally Economic indicators, Company-specific factors
@@
20090109 -- After a government report showed another big monthly drop in payrolls, resulting in the biggest annual job loss since just after World War II Economic indicators

### v3

@@ C Mathematical Proofs
where we ignore the subscript in (2.13) and denote $\widetilde{\mathbf{F}}^{s}$ as the estimated factor matrix from $\widetilde{\mathbf{X}}_{1}$ with sparsity $s$, and ${\mathbf{F}}^{s}$ as the true factor matrix retaining only keeps the top $s$ elements, both of which are normalized.
Proof. We only show it for $r=1$ since the proof is similar for $r>1$. Suppose the true factor ${\mathbf{F}}^{s}$ is the normalized true factor matrix such that ${\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s}/T={\mathbf{I}}_{r}$, which is similarly assumed in Bai and Ng (2013). By (2.13), we have
@@
Let Assumptions 1–7 hold. For any $s$ with $s<s_{0}$, there exists a constant $\tau_{s}>0$ such that
@@
Since each column in ${\mathbf{F}}^{s}({\mathbf{F}}^{s}{{}^{\prime}}{\mathbf{F}}^{s})^{-1}{\mathbf{F}}^{s}{{}^{\prime}}$ is a hat matrix and ${\mathbf{G}}\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}{\mathbf{G}}^{\prime}$ is a positive semi-definite matrix, then there exists a constant $\tau_{s}>0$ such that $I\geq\tau_{s}N_{2}T$. Then Lemma A.9 follows from this result and (S.25)–(C). This completes the proof. $\Box$
@@
Let Assumptions 1–7 hold. For any $s$ with $s\geq s_{0}$, we have
@@
$=$ $\mbox{tr}\{({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}-\widetilde{\mathbf{F}}^{s}(\widetilde{\mathbf{F}}^{s}{{}^{\prime}}\widetilde{\mathbf{F}}^{s})^{-1}\widetilde{\mathbf{F}}^{s}{{}^{\prime}}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s})\boldsymbol{\Lambda}_{2}^{\prime}\boldsymbol{\Lambda}_{2}({\mathbf{F}}^{0}-\widetilde{\mathbf{F}}^{s}){{}^{\prime}}\}.$ (S.32)
@@
By a similar argument for the remaining terms in (C), we can show that
@@
Proof of Theorem 4. We focus on the case where the number of partitions $J=1$, as the argument extends similarly to a general $J>1$. Under Assumption 3, which states that $s\asymp T$, we will show that for sufficiently large $T$,
$\lim_{N\rightarrow\infty}P(PC(s)<PC(s_{0}))=0,$
for all $s\neq s_{0}$ with $s\leq s_{\max}$. Since $s_{0}/T=c_{0}\in(0,1)$ by Assumption 3, for any $s$ such that $s/T=c_{s}\in(0,1)$, we adopt the theoretical convention in the proof that $s<s_{0}$ implies $c_{s}<c_{0}$, and similarly, $s>s_{0}$ implies $c_{s}>c_{0}$. We first consider the case when $s<s_{0}$. Note that
$PC(s)-PC(s_{0})=R(s,\widetilde{\mathbf{F}}^{s})-R(s_{0},\widetilde{\mathbf{F}}^{0})-r\frac{(s_{0}-s)}{T}C_{T}g(N_{1},T),$
@@
$P(PC(s)-PC(s_{0})<0)=P(\tau_{s}+O_{p}(C_{N_{1}T})<r\frac{(s_{0}-s)}{T}C_{T}g(N_{1},T))\rightarrow 0,$
since $C_{T}g(N_{1},T)\rightarrow 0$ and $C_{N_{1}T}$ is small enough for large $T$.
@@
$P(PC(s)-PC(s_{0})<0)=P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r\frac{(s-s_{0})}{T}C_{T}g(N_{1},T)).$
@@
Since $C_{N_{1}T}^{-2}C_{T}g(N_{1},T)\rightarrow\infty$, we have that
$P(R(s_{0},\widetilde{\mathbf{F}}^{0})-R(s,\widetilde{\mathbf{F}}^{s})>r\frac{(s-s_{0})}{T}C_{T}g(N_{1},T))\rightarrow 0,$
as $N\rightarrow\infty$ and $T$ is large enough. This completes the proof. $\Box$
D Description of the Identified Sparse Factors
@@
Table S.VII: Significant sparse factors over the time horizon from January 1, 2004, to December 31, 2016. Dates are formatted as yyyymmdd. The signs in the Factor Return column indicate whether the significant time factor resulted in a positive (++) or negative (-) return on that day. Reasons in the Reason column are extracted from daily reports on CNN Money (www.money.cnn.com) after the market closes each trading day. The Time Factor column summarizes the corresponding factor for each trading day with significant common risk. This website was shut down in late 2018 and its content was merged into the main CNN Business section.
@@
20070227 - A big decline in Chinese stocks, weakness in some key readings on the US economy and news that Vice President Dick Cheney was the apparent target in a Taliban suicide bombing attack in Afghanistan China, Economic indicators, Global events (terrorist attack)
@@
20070803 - Credit market fears, sparked by Wall Street bank Bear Stearns Credit risk, Market sentiment
20070809 - As new credit concerns surfaced, sparking a steep selloff Credit risk, Market sentiment
20070828 - Credit card defaults keep rising Credit risk
@@
20070918 ++ After the Federal Reserve cut a key short-term interest rate by a half-percentage point Government policies
20071019 - As Wachovia revives worries about bank sector Market sentiment
20071101 - On credit fears Credit risk, Market sentiment
20071107 - Renewed credit market fears Credit risk, Market sentiment
20071113 ++ After Wal-Mart’s earnings report Company-specific factors
@@
20071128 ++ Expectations the Federal Reserve will continue its rate cutting campaign and on strength from the embattled financials Government policies, Market sentiment
@@
20071211 - After the Federal Reserve cut the fed funds rate by a quarter-percentage point, as expected, but disappointed some investors looking for a bigger cut Government policies, Market sentiment
@@
20080104 - After a weaker-than-expected December jobs report exacerbated worries that the economy may be falling into recession Economic indicators, Market sentiment
20080115 F- After Citigroup’s weak earnings and a big drop in retail sales revived worries about the threat of recession Company-specific factors, Market sentiment
@@
20080117 - Recession worries following comments from Federal Reserve Chairman Ben Bernanke, Merrill Lynch’s big quarterly loss and weak readings on the housing and manufacturing sectors Market sentiment, Company-specific factors, Economic indicators
@@
20080123 ++ The Federal Reserve stepped in Tuesday and announced an emergency intermeeting interest rate cut, a decision that initially had a mixed impact on stocks, but helped pave the way for Wednesday’s bounce back Government policies
@@
20080205 - After a report showing a big slowdown in the services sector of the economy and cautionary comments from a Fed official amplified fears that a recession is underway or imminent Company-specific factors, Government policies
@@
20080229 - After AIG’s record loss added to worries about the financial sector and more weak economic news intensified fears about a recession Company-specific factors
20080306 - As investors eyed the latest wave of credit market woes Credit risk, Market sentiment
@@
20080311 ++ Announcement: the Federal Reserve will lend up to \$$200$ billion to banks and lenders as a means of loosening up tight credit markets Government policies
@@
20080318 ++ After the Federal Reserve cut the fed funds rate by three-quarters of a percentage point, surprising investors looking for a larger cut Government policies
@@
20080401 ++ Investors cheered signs that the companies hit hardest by the credit market crisis seem to be working through the problems Market sentiment
@@
20080606 - After oil prices spiked more than \$\\$ $11$ a barrel and the May jobs report showed a big jump in the unemployment rate Oil price, Economic indicators
20080626 - Selling accelerated following a record surge in oil prices Oil price
20080708 ++ Falling oil prices and a stronger dollar Oil price
@@
20080709 - More worries about Freddie Mac and Fannie Mae’s ability to raise capital exacerbated credit market and corporate profit jitters Market sentiment
@@
20080716 ++ Encouraging news from the banking and airline sectors Falling oil prices also helped spark a strong stock market rally Economic indicators, Oil price
@@
20080717 ++ A string of large oil price declines and more encouraging earnings results from the financial sector Oil price, Economic indicators
@@
20080724 - Renewed fears about the battered housing market and rising unemployment Market sentiment
@@
20080729 ++ Strong financial reports, rising consumer confidence and falling oil prices Economic indicators, Oil price
@@
20080805 ++ Oil prices fell sharply and investors appeared to take solace in the Federal Reserve’s assessment of the nation’s economy Oil price, Government policies
20080808 ++ Oil prices tumbled a 3-month low near \$\\$ $115$ a barrel Oil price
@@
20080904 - Mixed retail sales, lower oil prices and dour labor market readings amplified worries about a global economic slowdown Economic indicators, Oil price
@@
20080909 - Worries about Lehman Brothers’ ability to raise capital, and about the extent of AIG’s mortgage-related losses, exacerbated broad recession fears Company-specific factors
@@
20080915 - After one of the most calamitous days in US financial services history resulted in Bank of America’s \$$50$ billion deal to buy Merrill Lynch and the bankruptcy filing of Lehman Brothers Company-specific factors
@@
20080917 - The government’s emergency rescue of AIG amplified fears about the stability of financial markets Government policies
@@
20080918 ++ On a CNBC report that the government is working on a more permanent solution to absorbing bad debt Government policies
@@
20080919 ++ The government’s plan to help rescue banks from toxic mortgage debt soothed investors at the end of a gut-churning week on Wall Street Government policies
@@
20080922 - As investors worried about the specifics of the government’s \$$700$ billion bailout plan and rocketing oil prices - which saw its biggest one-day dollar gain ever Market sentiment
@@
20080929 - After the House rejected the government’s \$$700$ billion bank bailout plan Government policies
@@
20080930 ++ On bets that Congress will pass a version of the government’s \$$700$ billion package, following Monday’s crushing defeat Market sentiment, Government policies
@@
20081002 - As frozen credit markets and weak economic reports amplified jitters ahead of the House vote on the \$$700$ billion bank rescue plan Market sentiment, Government policies
@@
20081006 - As the \$$700$ billion bank bailout plan and European government attempts to prop up faltering banks failed to comfort panicky investors Market sentiment, Government policies
@@
20081007 - As the Federal Reserve’s plan to loosen credit markets failed to temper investor pessimism Market sentiment, Government policies
@@
20081009 - As panicked investors dumped stocks across the board Market sentiment
@@
20081013 ++ As investors bet that the worst of the credit crisis is over, following a series of global initiatives announced over the last few days Market sentiment
@@
20081015 - A weak retail sales report and dour forecasts from the Federal Reserve, coupled with sober comments from Fed Chairman Ben Bernanke, sent stocks tumbling Economic indicators, Government policies
@@
20081016 ++ As the lowest oil prices in more than a year gave investors a reason to scoop up shares battered in the recent market selloff Oil price
@@
20081020 ++ As investors welcomed talk of a second economic stimulus plan and an improvement in key lending rates Market sentiment, Government policies
@@
20081021 - As mixed corporate earnings reports gave investors a reason to retreat after the previous session’s big rally Economic indicators
@@
20081022 - As weak earnings and slumping oil prices amplified fears of a global recession Oil price
@@
20081024 - As Wall Street joined a worldwide market slump on bets that a recession is imminent - if not already under way Market sentiment
@@
20081027 - As recession jitters outweighed relief that the government’s programs to shore up the financial system have gotten underway Market sentiment
@@
20081028 ++ As investors dove back into stocks near the end of one of the worst months in Wall Street history Market sentiment
@@
20081031 ++ Capping off a strong week at the end of one of the worst months in Wall Street history Market sentiment
@@
20081104 ++ As millions of Americans battered by the weakened economy turned out to vote for the next President of the United States Market sentiment
@@
20081105 - As Barack Obama’s historic victory gave way to renewed worries about the struggling economy Market sentiment
@@
20081106 - As fears of a prolonged recession sent investors running for the exits Market sentiment
@@
20081110 - As ongoing recession fears overshadowed any relief about China’s \$$586$ billion stimulus plan and the government’s revamping of its deal with AIG China, Government policies
@@
20081111 - As recession fears trumped a new government and mortgage industry plan to help troubled homeowners Market sentiment, Government policies
@@
20081112 - As investors bet that a long and deep recession is on the horizon Market sentiment
@@
20081113 ++ As the major stock gauges bounced back from levels not seen since 2003 Market sentiment
@@
20081114 - As the worst retail sales on record ignited fears of a long recession Economic indicators
@@
20081117 - As investors eyed Citigroup’s massive job losses and a weak manufacturing report, while awaiting the fate of a potential bailout for the automakers Company-specific factors, Economic indicators
@@
20081119 - As ongoing anxiety about the economy and uncertainty about the future of the auto industry weighed on the market Market sentiment
@@
20081120 - As fears of a prolonged recession sparked a massive selloff Market sentiment
@@
20081121 ++ After reports surfaced that President-elect Barack Obama will nominate New York Federal Bank President Timothy Geithner as his new Treasury secretary Government policies
@@
20081124 ++ As Citigroup’s massive rescue package and President-elect Obama’s picks for his economic team pushed investors off the sidelines Company-specific factors, Government policies
@@
20081126 ++ As investors scooped up stocks hit in the recent selloff, ahead of the Thanksgiving holiday Market sentiment
@@
20081201 - As investors bailed out following confirmation that the US is mired in a recession and indications that it’s likely to continue for some time Market sentiment
@@
20081202 ++ As investors welcomed signs that the automakers might get a bailout after all Government policies
20081203 ++ Oil prices fall after government inventory report Oil price
@@
20081204 - As a rash of job cuts at major companies added to jitters ahead of the November jobs report Economic indicators
@@
20081205 ++ After a brutal November employment report, as investors extended the recent trend of buying despite the bad news Economic indicators
@@
20081208 ++ As investors welcomed President-elect Barack Obama’s plan to create jobs and revive the economy, and reports that government help for the automakers is on the way Government policies
@@
20081211 - On worries that the \$\\$ $14$ billion auto rescue bill won’t pass in the Senate due to Republican opposition Government policies
@@
20081216 ++ After the Federal Reserve cut a key short-term interest rate to the lowest level on record, and signaled it had more tools available to help the economy as the recession stretches on Government policies, Market sentiment
@@
20081222 - Amid concerns about fourth-quarter corporate earnings, falling oil prices and ongoing woes in the auto industry Market sentiment, Oil price, Economic indicators
@@
20081230 ++ As investors scooped up a variety of shares hit hard in the 2008 stock market battering Market sentiment
@@
20090102 ++ With investors starting off a new year on the right foot, after an abysmal 2008 Market sentiment
@@
20090106 ++ As investors looked beyond the Federal Reserve’s dour outlook on the economy and instead scooped up shares hit in last year’s big selloff Market sentiment
@@
20090107 - After weak labor market reports and dour forecasts from Alcoa and Intel gave investors reasons to retreat after the recent rally Economic indicators, Company-specific factors
@@
20090109 - After a government report showed another big monthly drop in payrolls, resulting in the biggest annual job loss since just after World War II Economic indicators